Slope = 1.80. Intercept 41.50 Question 10 x Points: o out of i Suppose that the regression for predicting weight (in pounds) from Height (in inches) is given by Weight -115+3.6(Height) Which of the following statements is correct? I. A person who is 61 inches tall will weigh 104.6 poulds. II. For every additional inch of height, the predicted weight will increase, on average by 3.6 pounds III. The correlation between weight and height is negative. A. II only B. Iand II only C. I only D. II and III only E. III only Question 11 search

Need solution of this question with explanation!
Answer #1

Solution:- (B) I and II only.

I) A person who is 61 inches tall will weigh 104.6 pounds.

Weight = -115 + 3.6*Height

Weight = - 115 + 3.6*61

Weight = 104.6

Hence statement I is true.

II) The slope of regression line is 3.6, hence for every additional inch of height, the predicted weight will increase, on average by 3.6 pounds.

III)

Weight = -115 + 3.6*Height

Since the slope of the regression line (3.6) is positive, hence the correlation between weight and height would also be positive.

Hence statement III is not correct.